# Reflection of Nichols Algebras over Coquasi-Hopf Algebras

### v1

@@ 5.1 Definition of reflection and basic properties
Such a question arises naturally: how do the irreducibility properties behave under reflections? The following result provides a reassuring answer.
@@
We want to prove $\langle L\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. To see this, take any homogeneous $A\in\mathcal{A}(N)$. Then for all homogeneous $X\in L$, by Lemma 2.6,
@@
$=\langle f,s(\mathcal{S}(A_{1}),(\operatorname{ad}(A_{4})(X_{0}))_{-1}A_{6})t(A_{3},X_{-1})\mathcal{S}(A_{2})(\operatorname{ad}(A_{4})(X_{0}))_{-2}A_{5})\rangle(\operatorname{ad}(A_{4})(X_{0}))_{0}.$
@@
Hence $\operatorname{ad}(A)(\langle f,X_{-1}\rangle X_{0})\in\langle L\rangle.$ This ensures the first axiom of a Yetter-Drinfeld module, and the third axiom holds automatically. Since $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ is generated by $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ as a Yetter-Drinfeld module and is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 4.8 (2). Then $L$ must generate the entire space. This forces $L=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}),$ establishing the irreducibility.
@@
Suppose $M\in\mathcal{F}_{\theta}$ and $M$ admits $i$-th reflection for each $i\in\mathbb{I}$. We define $a_{ii}^{M}=2$ for all $1\leq i\leq\theta$ and define $a_{ij}^{M}=-m_{ij}^{M}$. Then $(a_{ij}^{M})_{i,j\in\mathbb{I}}$ is a generalized Cartan matrix.
@@
Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so is $N$. Furthermore, $R_{i}(M)\cong R_{i}(N)$ and $a_{ij}^{M}=a_{ij}^{N}$ for each $j\in\mathbb{I}$.
@@
We denote $\phi:M\rightarrow N$ be isomorphism of tuples with each component $\phi_{j}:M_{j}\rightarrow N_{j}$ an isomorphism of Yetter-Drinfeld modules. We first demonstrate that $\phi_{i}$ inducing a unique morphism $\mathcal{B}(\phi_{i}):\mathcal{B}(M_{i})\rightarrow\mathcal{B}(N_{i})$.
@@ 5.2 Main Theorem
In this subsection, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ is irredcuible, $i\in\mathbb{I}$. Furhtermore, we assume $\mathcal{B}(M_{i})$ is finite-dimensional for each $i\in\mathbb{I}$. By Corollary 3.8, we have such a braided tensor equivalence for each $i\in\mathbb{I}$.
@@
The main result of this paper proves that the reflection of Nichols algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$} gives rise to a semi-Cartan graph. The first lemma provides some relations between reflection of $M$ and the braided tensor equivalence $\Omega_{i}$.
@@
The next theorem gives a natural explanation of reflections of tuples of Yetter-Drinfeld modules. Although the proof strategy parallels the Hopf algebra case [rootsys, Theorem 13.4.9], we include the details here for the sake of completeness and the reader’s convenience.
@@
With above assumptions on $M$. For each $i\in\mathbb{I}$, suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
@@ 6 Applications
Building upon the framework developed in the preceding sections, we now present some applications of the semi-Cartan graph theory for Nichols algebras over certain coquasi-Hopf algebras. We establish several criteria that connect the finite-dimensionality of a Nichols algebra to the finiteness and structure of its associated semi-Cartan graph and Weyl groupoid. Then we apply this theory to provide a new proof for the key result in [huang2024classification].
@@ 6.1 A criterion for finite-dimensional Nichols algebras
If we regard $\mathcal{B}(M)$ as an $\mathbb{N}_{0}$-graded object in $^{G}_{G}\mathcal{YD}^{\Phi}$}. By [defofnichols, Lemma 2.6] the graded dual of $\mathcal{B}(M)$ is isomorphic to $\mathcal{B}(M^{*})$. We thus obtain an isomorphism of $\mathbb{N}_{0}$-graded objects in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
@@
Since $A\otimes B\cong B\otimes A$ as $\mathbb{Z}^{\theta}$-graded vector spaces. the supports on both sides coincide, which establishes the first claim.
@@
is called the set of real roots of $\mathcal{G}$ at $X$. We call the semi-Cartan graph $\mathcal{G}$ is finite, if $\Delta^{X,\operatorname{re}}$ is finite for all $X\in\mathcal{X}$.
@@
We recall that every element $\omega\in\operatorname{Hom}(\mathcal{W}(\mathcal{G}),M)$ can be expressed as an iteration of simple reflections. It follows that Thus $\Delta^{M\operatorname{re}}\subseteq\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*})$. Now, since $\mathcal{B}(M)$ is finite-dimensional, the set $\operatorname{Supp}\mathcal{B}(M)$ is finite. Moreover, as vector spaces, $\mathcal{B}(M)\cong\mathcal{B}(M^{*})^{\mathrm{gr\text{-}dual}}$, so $\operatorname{Supp}\mathcal{B}(M^{*})$ is also finite. Combining these observations, we conclude that $\Delta^{M,\operatorname{re}}$ is finite. The same reasoning applies to any $N\in X$: we have
@@
[HS10, Lemma 5.1] Suppose $\mathcal{G}=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$ is a semi-Cartan graph. Then $\mathcal{G}$ is finite semi-Cartan graph if and only if the Weyl groupoid $\mathcal{W}(\mathcal{G})$ is finite.
@@
The proof of next theorem is similar to the Hopf algebra situations, we write down here for completeness.
@@ 6.2 A class of infinite-dimensional Nichols algebras
The purpose of this section is to give a new proof of Proposition 4.1 in huang2024classification through applying our previous observations.
@@
(ii) $\mathcal{B}(V_{1})$, $\mathcal{B}(V_{2})$, $\mathcal{B}(V_{3})$ are finite-dimensional Nichols algebras by [huang2024classification, Proposition 2.12]. Moreover, $\mathcal{B}(V_{1}\oplus V_{2})$, $\mathcal{B}(V_{1}\oplus V_{3})$, $\mathcal{B}(V_{2}\oplus V_{3})$ are finite-dimensional by [nondiagonal Proposition 5.1].
@@
[huang2024classification, Proposition 4.1] With above notations, $\mathcal{B}(V_{1}\oplus V_{2}\oplus V_{3})$ is infinite-dimensional.
It should be emphasized that the above theorem plays the key role in that paper and the original proof relays on heavy computations. The purpose of the following context is to give a new proof of this theorem through applying reflection theory of Nichols algebras.
@@
be a $3$-cocycle on $H$, where $0\leq i_{l},j_{l},k_{l}\leq 1$, $1\leq l\leq 3$. By huang2024classification Lemma 3.5, we may give a complete list of irreducible Yetter-Drinfeld modules. In particular, restricting to those irreducible Yetter–Drinfeld modules that generate finite-dimensional Nichols algebras, the isomorphism classes are given by the set
@@
The category ${{}^{H}_{H}\mathcal{YD}^{\phi}}_{\operatorname{fd}}$ is rigid. It is straightforward to verify that for each $1\leq i\leq 6$, $W_{i}^{*}$ does exist with $W_{i}^{*}\cong W_{i}$. This implies $W_{i}^{*}\in S$ as well. Moreover, $\mathcal{B}(W_{i})\cong\mathcal{B}(W_{i}^{*})$ as Hopf algebras in $^{H}_{H}\mathcal{YD}$}.
@@
This proof involves extensive but analogous computations. We begin by establishing that $\operatorname{ad}(W_{1})^{2}(W_{2})=0.$ To this end, we compute the coproducts:
@@
This combines (6.8) implies $\operatorname{ad}(W_{1})^{2}(W_{3})=0$. As in the first case, we can derive the following identities:
@@
All cases are leading to contradictions. This completes the induction. ∎
@@
By [QQG, Lemma 4.4], we deduce that
@@
in the sense of [QQG, Definition 4.3]. It follows that $\mathcal{B}(V)$ is infinite-dimensional. ∎
Remark 6.15.
For the sake of simplicity, this article studies only the cosemisimple pointed coquasi-Hopf algebras under minor conditions. We hope to extend this theory to all coquasi-Hopf algebras with bijective antipodes.

### v2

@@ 5.1 Definition of reflection and basic properties
A natural question arises: how do the irreducibility properties behave under reflections? The following result provides a reassuring answer.
@@
We want to prove $\langle L\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. To see this, take any homogeneous $A\in\mathcal{A}(M_{i})$. Then for all homogeneous $X\in L$, by Lemma 2.6,
@@
$=\langle f,s(\mathcal{S}(A_{1}),(\operatorname{ad}(A_{4})(X_{0}))_{-1}A_{6})t(A_{3},X_{-1})\mathcal{S}(A_{2})((\operatorname{ad}(A_{4})(X_{0}))_{-2}A_{5})\rangle(\operatorname{ad}(A_{4})(X_{0}))_{0}.$
@@
Hence $\operatorname{ad}(A)(\langle f,X_{-1}\rangle X_{0})\in\langle L\rangle.$ This ensures the first axiom of a Yetter-Drinfeld module, and the third axiom holds automatically. Therefore $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}.
Since $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 4.8 (2), and $P\neq 0$, we have $\langle P\rangle=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$. It is direct to see that $\langle P\rangle=\bigoplus_{i=0}^{m_{ij}}\langle P\rangle\cap\mathcal{B}^{i}(M_{i}\oplus M_{j})$. Then
$P=\langle P\rangle\cap\mathcal{B}^{m_{ij}}(M_{i}\oplus M_{j})=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})\cap\mathcal{B}^{m_{ij}}(M_{i}\oplus M_{j})=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}).$
This establishes the irreducibility. ∎
@@
Suppose $M\in\mathcal{F}_{\theta}$ and $M$ admits the $i$-th reflection for each $i\in\mathbb{I}$. We define $a_{ii}^{M}=2$ for all $1\leq i\leq\theta$ and define $a_{ij}^{M}=-m_{ij}^{M}$. Then $(a_{ij}^{M})_{i,j\in\mathbb{I}}$ is a generalized Cartan matrix.
@@
Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so does $N$. Furthermore, $R_{i}(M)\cong R_{i}(N)$ and $a_{ij}^{M}=a_{ij}^{N}$ for each $j\in\mathbb{I}$.
@@
Let $\phi:M\rightarrow N$ be an isomorphism of tuples with each component $\phi_{j}:M_{j}\rightarrow N_{j}$ an isomorphism of Yetter-Drinfeld modules. We first demonstrate that $\phi_{i}$ inducing a unique morphism $\mathcal{B}(\phi_{i}):\mathcal{B}(M_{i})\rightarrow\mathcal{B}(N_{i})$.
@@ 5.2 Main Theorem
In this subsection, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ being irreducible, $i\in\mathbb{I}$. Furthermore, we assume $\mathcal{B}(M_{i})$ is finite-dimensional for each $i\in\mathbb{I}$. By Corollary 3.8, we have such a braided tensor equivalence for each $i\in\mathbb{I}$.
@@
The main result of this paper proves that the reflection of Nichols algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$} gives rise to a semi-Cartan graph. The first lemma provides some relations between reflection of $M$ and the braided tensor equivalence $\Omega_{i}$.
@@
The next theorem gives a natural explanation of reflections of tuples of Yetter-Drinfeld modules. Although the proof strategy parallels that of the Hopf algebra case [Heckenberger and Schneider, [2020] ©2020, Theorem 13.4.9], we include the details here for the sake of completeness and the reader’s convenience.
@@
With the above assumptions on $M$. For each $i\in\mathbb{I}$, suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
@@ 6 Applications
Building upon the framework developed in the preceding sections, we now present some applications of the semi-Cartan graph theory to Nichols algebras over certain coquasi-Hopf algebras. We establish several criteria that connect the finite-dimensionality of a Nichols algebra to the finiteness and structure of its associated semi-Cartan graph and Weyl groupoid. Then we apply this theory to provide a new proof for the key result in [Huang et al., ].
@@ 6.1 A criterion for finite-dimensional Nichols algebras
If we regard $\mathcal{B}(M)$ as an $\mathbb{N}_{0}$-graded object in $^{G}_{G}\mathcal{YD}^{\Phi}$}, by [Meir, 2022, Lemma 2.6] the graded dual of $\mathcal{B}(M)$ is isomorphic to $\mathcal{B}(M^{*})$. We thus obtain an isomorphism of $\mathbb{N}_{0}$-graded objects in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
@@
Since $A\otimes B\cong B\otimes A$ as $\mathbb{Z}^{\theta}$-graded vector spaces, the supports on both sides coincide, which establishes the first claim.
@@
is called the set of real roots of $\mathcal{G}$ at $X$. We call the semi-Cartan graph $\mathcal{G}$ finite, if $\Delta^{X,\operatorname{re}}$ is finite for all $X\in\mathcal{X}$.
@@
We recall that every element $\omega\in\operatorname{Hom}(\mathcal{W}(\mathcal{G}),M)$ can be expressed as an iteration of simple reflections. It follows that $\Delta^{M\operatorname{re}}\subseteq\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*})$. Now, since $\mathcal{B}(M)$ is finite-dimensional, the set $\operatorname{Supp}\mathcal{B}(M)$ is finite. Moreover, as vector spaces, $\mathcal{B}(M)\cong\mathcal{B}(M^{*})^{\mathrm{gr\text{-}dual}}$, so $\operatorname{Supp}\mathcal{B}(M^{*})$ is also finite. Combining these observations, we conclude that $\Delta^{M,\operatorname{re}}$ is finite. The same reasoning applies to any $N\in X$: we have
@@
[Heckenberger and Schneider, 2010b, Lemma 5.1] Suppose $\mathcal{G}=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$ is a semi-Cartan graph. Then $\mathcal{G}$ is finite semi-Cartan graph if and only if the Weyl groupoid $\mathcal{W}(\mathcal{G})$ is finite.
@@
The proof of the next theorem is similar to the Hopf algebra situation; we include it here for completeness.
@@ 6.2 A class of infinite-dimensional Nichols algebras
The purpose of this section is to give a new proof of Proposition 4.1 in Huang et al. through applying our previous observations.
@@
(ii) $\mathcal{B}(V_{1})$, $\mathcal{B}(V_{2})$, $\mathcal{B}(V_{3})$ are finite-dimensional Nichols algebras by [Huang et al., , Proposition 2.12]. Moreover, $\mathcal{B}(V_{1}\oplus V_{2})$, $\mathcal{B}(V_{1}\oplus V_{3})$, $\mathcal{B}(V_{2}\oplus V_{3})$ are finite-dimensional by [Huang et al., 2018, Proposition 5.1].
@@
[Huang et al., , Proposition 4.1] With above notations, $\mathcal{B}(V_{1}\oplus V_{2}\oplus V_{3})$ is infinite-dimensional.
It should be emphasized that the above theorem plays a key role in that paper and the original proof relies on heavy computations The purpose of the following discussion is to give a new proof of this theorem by applying the reflection theory of Nichols algebras.
@@
be a $3$-cocycle on $H$, where $0\leq i_{l},j_{l},k_{l}\leq 1$, $1\leq l\leq 3$. By [Huang et al., , Lemma 3.5], we may give a complete list of irreducible Yetter-Drinfeld modules. In particular, restricting to those irreducible Yetter–Drinfeld modules that generate finite-dimensional Nichols algebras, the isomorphism classes are given by the set
@@
The category ${{}^{H}_{H}\mathcal{YD}^{\phi}}_{\operatorname{fd}}$ is rigid. It is straightforward to verify that for each $1\leq i\leq 6$, $W_{i}^{*}$ does exist with $W_{i}^{*}\cong W_{i}$. This implies $W_{i}^{*}\in S$ as well. Moreover, $\mathcal{B}(W_{i})\cong\mathcal{B}(W_{i}^{*})$ as Hopf algebras in $^{H}_{H}\mathcal{YD}^{\phi}$}.
@@
This proof involves extensive but analogous computations.
Step1
We begin by establishing that $\operatorname{ad}(W_{1})^{2}(W_{2})=0.$ To this end, we compute the coproducts:
@@
This argument, combined with (6.8), implies $\operatorname{ad}(W_{1})^{2}(W_{3})=0$. As in the first case, we can derive the following identities:
@@
All of these cases lead to contradictions. This completes the induction. ∎
@@
By [Huang et al., $20$, Lemma 4.4], we deduce that
@@
in the sense of [Huang et al., $20$, Definition 4.3]. It follows that $\mathcal{B}(V)$ is infinite-dimensional. ∎